### v1

# A Compact Post-quantum Strong Designated Verifier Signature Scheme from Isogenies

Abstract
Digital signatures are essential cryptographic tools that provide authentication and integrity in digital communications. However, privacy-sensitive applications—such as e-voting and digital cash—require more restrictive verification models to ensure confidentiality and control. Strong Designated Verifier Signature (SDVS) schemes address this need by enabling the signer to designate a specific verifier, ensuring that only this party can validate the signature. Existing SDVS constructions are primarily based on number-theoretic assumptions and are therefore vulnerable to quantum attacks. Although post-quantum alternatives—particularly those based on lattices—have been proposed, they often entail large key and signature sizes.
In this work, we introduce CSI-SDVS, a novel isogeny-based SDVS scheme that offers a compact, quantum-resistant alternative. Our construction builds on the ideal class group action framework of CSIDH and the signature techniques of CSI-FiSh, and relies on the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP). CSI-SDVS achieves strong security guarantees—namely, Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI)—in the random oracle model. Remarkably, both the keys and signatures in CSI-SDVS are of size $\mathcal{O}(\lambda)$, representing a significant improvement over the typical $\mathcal{O}(\lambda^{2})$ bounds in existing post-quantum SDVS schemes, thereby making it among the most compact PQC-based SDVS schemes and the only post-quantum secure construction based on isogenies.
Keywords: Digital Signature; Strong Designated Verifier; Post-quantum Cryptography; Isogeny-based Cryptography
1 Introduction
In many cryptographic applications, digital signatures play a central role by ensuring message authenticity and protecting against tampering. Yet, their default public verifiability model may not align with scenarios that demand selective disclosure or restricted verification. In particular, contexts involving sensitive or confidential data—such as personalized medical documents, legal contracts, or anonymous transactions—necessitate a finer level of control over who can verify the legitimacy of a signature. These limitations have motivated the exploration of alternative signature paradigms that can embed verification constraints directly into the signing process.
To address these concerns, Jakobsson et al. [1] and Chaum [2] independently introduced the notion of Designated Verifier Signatures (DVS) in 1996. In a DVS scheme, only a specific verifier is capable of validating the authenticity of a signature, while any third party remains unconvinced. This is achieved by enabling the designated verifier to simulate signatures that are computationally indistinguishable from authentic ones, thereby preserving ambiguity about their origin. Even if a valid-looking signature is disclosed, only the designated verifier can be confident of its validity. This crucial feature—referred to as non-transferability—ensures that the verification outcome is non-provable to others.
Building upon this idea, the concept of Strong Designated Verifier Signatures (SDVS) later emerged as an extension of the standard DVS framework, offering enhanced privacy guarantees. Although briefly mentioned in the appendix of Jakobsson et al.’s foundational work [1], a formal definition and construction were first provided by Laguillaumie et al. [3] in 2004. The key advancement in SDVS lies in the protection of the signer’s identity against any third party, including passive observers, even in the presence of multiple potential signers. In particular, SDVS schemes satisfy the Privacy of Signer’s Identity property: given a valid signature and two candidate signing public keys, it is computationally infeasible for anyone other than the designated verifier to determine which key was used. This is accomplished by incorporating the designated verifier’s secret key into the verification algorithm, rendering the process inherently non-transferable and private.
1.1 Related Work
Strong designated verifier signature schemes have been extensively studied in both classical and post-quantum cryptographic settings.
Classical Constructions. Independent of the foundational work by Laguillaumie et al. [3], Saeednia et al. [4] proposed an efficient SDVS scheme by adapting the Schnorr signature and incorporating the designated verifier’s secret key into the verification process. Subsequently, Huang et al. [5] and Kang et al. [6] introduced identity-based SDVS (IBSDVS) schemes and proved their unforgeability under the Bilinear Diffie–Hellman (BDH) assumption. However, most classical SDVS schemes rely on number-theoretic hardness assumptions—such as integer factorization and discrete logarithms—which are rendered insecure in the presence of quantum adversaries, as Shor’s algorithm [7] can efficiently solve these problems.
Post-Quantum Constructions. In response to the vulnerabilities of classical assumptions, post-quantum SDVS schemes have been developed based on quantum-resistant foundations. Code-based SDVS constructions were proposed by Assar [8], Ren et al. [9], and Shooshtari et al. [10], leveraging the hardness of the bounded syndrome decoding problem in coding theory [11]. While these schemes are believed to be quantum-resistant, it was shown in [12] that the constructions by Ren et al. and Shooshtari et al. fail to satisfy the critical non-transferability property, which is central to the SDVS notion.
One of the earliest lattice-based SDVS constructions was introduced by Wang et al. [14], employing preimage sampleable functions (PSFs) and Bonsai Trees for trapdoor delegation. Their work builds on the influential framework of Gentry et al. [13], which introduced PSFs as a core primitive in lattice-based signatures. Although the construction offers provable security in the random oracle model, it suffers from efficiency issues due to large key and signature sizes. Later, in 2016, Noh et al. [15] proposed a scheme in the standard model based on an LWE-based encryption scheme and a lattice-based chameleon hash function, albeit at the cost of significant computational overhead. To improve practicality, Cai et al. [16] introduced a more efficient scheme based on the Ring-SIS assumption, using rejection sampling to reduce signature size while maintaining security. More recently, Zhang et al. [17] proposed an SDVS construction combining SIS and LWE assumptions, further contributing to the development of quantum-secure SDVS protocols. However, lattice-based schemes generally incur key and signature sizes with complexity $\mathcal{O}(\lambda^{2})$ [14, 15, 16, 17], which may limit their suitability for resource-constrained environments.
It is worth mentioning that Sun et al. [18] proposed an isogeny-based SDVS scheme built upon the Supersingular Isogeny Diffie–Hellman (SIDH) protocol [19]. However, this construction is no longer considered secure due to several successful cryptanalytic attacks [20, 21, 22] against SIDH. Furthermore, the scheme employed deterministic signing, which may introduce privacy and replayability concerns in certain scenarios.
1.2 Our Contribution
Our main contributions are as follows:
  • We present CSI-SDVS, an efficient post-quantum SDVS scheme constructed via class group actions on the set of $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves. The security of our scheme is based on the hardness of the Multi-Target Group Action Inverse Problem (MT-GAIP) [24], and the design draws inspiration from the CSIDH [23] and CSI-FiSh [24] frameworks.
  • The scheme provides strong security guarantees, including Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI), making it well-suited for applications requiring strict confidentiality and verifier-specific authentication.
  • In terms of efficiency, our construction achieves key and signature sizes of asymptotic complexity $\mathcal{O}(\lambda)$, significantly improving over the $\mathcal{O}(\lambda^{2})$ bounds common in existing lattice-based SDVS schemes [14, 15, 16, 17].
  • To the best of our knowledge, this is the only SDVS scheme based on isogenies that offers post-quantum security, contributing a new design within the scope of SDVS schemes.
Outline.
In the next section, we recall the formal definition of an SDVS scheme and provide the associated security requirements that such a scheme must satisfy. We then present a brief overview of the necessary mathematical prerequisites and describe the underlying group action-based hardness assumptions upon which our construction relies. In Section 3, we introduce our proposed SDVS scheme, CSI-SDVS. Section 4 analyzes the efficiency of our construction, and Section 5 provides a formal security proof.
2 Preliminaries
Notation. We denote the security parameter by $\lambda$. For a finite set $\mathbb{X}$, the notation $x\xleftarrow{\$}\mathbb{X}$ indicates that $x$ is sampled uniformly at random from $\mathbb{X}$. A probabilistic polynomial-time ($\mathsf{PPT}$) algorithm $A$ that outputs a value $x$ on input $y$ is denoted by $x\xleftarrow{\$}A(y)$. In contrast, for a deterministic polynomial-time ($\mathsf{DPT}$) algorithm, we write $x:=A(y)$. A function $\epsilon:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$ is said to be negligible if for every positive polynomial $p(\cdot)$, there exists a threshold $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$, it holds that $\epsilon(n)<\frac{1}{p(n)}$. The concatenation of two bit strings $s_{1}$ and $s_{2}$ is written as $s_{1}\|s_{2}$. For any positive integer $n$, we define $\mathbb{Z}_{n}:=\mathbb{Z}/n\mathbb{Z}$. We use the notation $\forall i\in[T]$ to denote ”for all $i\in\{1,2,\ldots,T\}$”.
2.1 Strong Designated Verifier Signature (SDVS) Scheme
An SDVS scheme involves two distinct parties: a signer $\mathsf{S}$ and a designated verifier $\mathsf{V}$, and is defined over a message space $\mathcal{M}$ and a signature space $\mathsf{\Sigma}$. In such a scheme, the signer $\mathsf{S}$, holding a secret key $\mathsf{sk_{S}}$, generates a signature $\sigma\in\mathsf{\Sigma}$ on a message $m\in\mathcal{M}$ using its secret key and the public key of the designated verifier. The resulting signature $\sigma$ can be verified only by $\mathsf{V}$ using its secret key $\mathsf{sk_{V}}$, ensuring that no third party can verify the signature or be convinced of its validity. This restriction preserves the signer’s privacy by eliminating public verifiability while maintaining correctness for the designated verifier. Moreover, the designated verifier is equipped with a simulation algorithm $\mathsf{Simul}$ that can generate a simulated signature $\sigma’$ ’ on any message $m’\in\mathcal{M}$ using $\mathsf{sk_{V}}$ and the signer’s public key, thereby preventing any verifiable evidence of the signer’s involvement. The formal syntax of an SDVS scheme is presented below, following the framework introduced in [18].
Definition 2.1 (SDVS Scheme).
A strong designated verifier signature scheme is defined as a tuple of six polynomial-time algorithms:
$\mathsf{Setup}(1^{\lambda})\xrightarrow{\$}\mathsf{pp}$:
A PPT algorithm that, given a security parameter $\lambda$, outputs the public parameters $\mathsf{pp}$.
$\mathsf{SigKeyGen}(\mathsf{pp})\xrightarrow{\$}(\mathsf{sk_{S}},\mathsf{pk_{S}})$:
A PPT algorithm that, on input $\mathsf{pp}$, generates a secret/public key pair $(\mathsf{sk_{S}},\mathsf{pk_{S}})$ for the signer S.
$\mathsf{VerKeyGen}(\mathsf{pp})\xrightarrow{\$}(\mathsf{sk_{V}},\mathsf{pk_{V}})$:
A PPT algorithm that, on input $\mathsf{pp}$, generates a secret/public key pair $(\mathsf{sk_{V}},\mathsf{pk_{V}})$ for the designated verifier V.
$\mathsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m)\xrightarrow{\$}\sigma$:
A PPT (or DPT) algorithm that, given the public parameters $\mathsf{pp}$, the signer’s secret key $\mathsf{sk_{S}}$, the public key $\mathsf{pk_{V}}$ of the designated verifier, and a message $m\in\mathcal{M}$, outputs a real designated verifier signature $\sigma\in\mathsf{\Sigma}$.
$\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m,\sigma)\rightarrow 0/1$:
A DPT algorithm that, given the public parameters $\mathsf{pp}$, the designated verifier’s secret key $\mathsf{sk_{V}}$, the signer’s public key $\mathsf{pk_{S}}$, a message $m\in\mathcal{M}$, and a signature $\sigma\in\mathsf{\Sigma}$, outputs a decision bit $b\in\{0,1\}$, where $b=1$ indicates acceptance and $b=0$ indicates rejection.
$\mathsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m)\xrightarrow{\$}\sigma^{\prime}$:
A PPT (or DPT) algorithm that, given the public parameters $\mathsf{pp}$, the designated verifier’s secret key $\mathsf{sk_{V}}$, the signer’s public key $\mathsf{pk_{S}}$, and a message $m\in\mathcal{M}$, outputs a simulated designated verifier signature $\sigma^{\prime}\in\mathsf{\Sigma}$.
Correctness. Let $\mathsf{pp}$ be the public parameters generated by $\mathsf{Setup}(1^{\lambda})$, and let $(\mathsf{sk_{S}},\mathsf{pk_{S}})$ and $(\mathsf{sk_{V}},\mathsf{pk_{V}})$ be the key pairs generated by $\mathsf{SigKeyGen}(\mathsf{pp})$ and $\mathsf{VerKeyGen}(\mathsf{pp})$, respectively. The correctness of an SDVS scheme requires that, for any message $m\in\mathcal{M}$, the following conditions hold with overwhelming probability:
$\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m,\mathsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m))=1,$
for a real designated verifier signature, and
$\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m,\mathsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m))=1,$
for a simulated designated verifier signature. These conditions ensure that honestly generated real and simulated designated verifier signatures are both valid and accepted by the verifier.
2.2 Security Model
An SDVS scheme must ensure Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI). The formal security definitions corresponding to these properties are presented below.
Strong Unforgeability under Chosen-Message Attacks (SUF-CMA) Security
In our SDVS scheme, CSI-SDVS, presented in Section 3, we adopt the notion of strong unforgeability, which is a stricter security requirement than existential unforgeability (EUF-CMA). This notion guarantees that no PPT adversary—without access to the secret key of either the signer or the designated verifier—can generate a valid designated verifier signature, even on messages that were previously queried via either the signing or simulating oracles. The formal definition is given below.
Definition 2.2 (SUF-CMA Security).
A strong designated verifier signature (SDVS) scheme is strongly unforgeable under chosen message attacks (SUF-CMA) if the advantage of any PPT adversary $\mathcal{A}$, defined as
$\mathsf{Adv}^{\mathsf{SUF\text{-}CMA}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)=\Pr\left[\mathcal{A}\text{ wins }\mathsf{Exp}^{\mathsf{SUF\text{-}CMA}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)\right],$
is negligible. Here, $\mathsf{Exp}^{\mathsf{SUF\text{-}CMA}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)$ denotes the SUF-CMA experiment between a challenger $\mathcal{C}$ and the adversary $\mathcal{A}$, as illustrated in Figure 1.
Setup: The challenger $\mathcal{C}$ runs $\textsf{Setup}(1^{\lambda})$ to get $\mathsf{pp}$, $\textsf{SigKeyGen}(\mathsf{pp})$, and $\textsf{VerKeyGen}(\mathsf{pp})$ to get $(\mathsf{sk_{S}},\mathsf{pk_{S}})$ and $(\mathsf{sk_{V}},\mathsf{pk_{V}})$ for the signer $\mathsf{S}$ and the designated verifier $\mathsf{V}$, respectively. $\mathcal{C}$ keeps $(\mathsf{sk_{S}},\mathsf{sk_{V}})$ in secret and sends $(\mathsf{pp},\mathsf{pk_{S}},\mathsf{pk_{V}})$ to the adversary $\mathcal{A}$. It also sets the list of message-signature pairs $\mathcal{Q}$ to $\varnothing$.
Query Phase: In this phase, $\mathcal{C}$ responds to polynomially many adaptive queries made by $\mathcal{A}$ by following the steps described below: Signing Oracle $\mathcal{O}_{\mathsf{Sign}}$: On receiving queries on a message $m$ from $\mathcal{A}$, $\mathcal{C}$ runs $\textsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m)$ to obtain a signature $\sigma$ and sends it to $\mathcal{A}$. It then sets $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$. Simulating Oracle $\mathcal{O}_{\mathsf{Simul}}$: On receiving queries on a message $m$ from $\mathcal{A}$, $\mathcal{C}$ runs $\textsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m)$ to obtain a signature $\sigma$ and sends it to $\mathcal{A}$. It then sets $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$. Verifying Oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving a message-signature pair $(m,\sigma)$ adaptively from $\mathcal{A}$, $\mathcal{C}$ runs $\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},\sigma,m)$ to obtain a decisional value $1$ for valid and 00 otherwise, and sends it to $\mathcal{A}$. Output: Eventually, $\mathcal{A}$ outputs a forgery $(m^{\ast},\sigma^{\ast})$, and wins if the following conditions hold: $\textsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},\sigma^{\ast},m^{\ast})=1.$ $(m^{\ast},\sigma^{\ast})\not\in\mathcal{Q}.$
Figure 1: $\mathsf{Exp}_{\mathsf{SDVS},\mathcal{A}}^{\mathsf{SUF\text{-}CMA}}$ Experiment.
The probability is taken over the randomness of $\mathcal{A}$, and the algorithms $\mathsf{Sign}$ and $\mathsf{Simul}$.
Non-Transferability (NT) Security
In an SDVS scheme, the notion of non-transferability (NT) ensures that the designated verifier cannot convince any third party of the authenticity of a valid signature. This is achieved by enabling the verifier to use the Simul algorithm to generate simulated signatures that are computationally indistinguishable from genuine ones produced by the signer. As a result, any conviction about the validity of a signature cannot be reliably transferred to others. The NT security of our SDVS construction is formalized as follows:
Definition 2.3 (NT Security).
A strong designated verifier signature (SDVS) scheme satisfies non-transferability (NT) security if the advantage of any PPT adversary $\mathcal{A}$, defined as
$\mathsf{Adv}^{\mathsf{NT}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)=\Pr\left[\mathcal{A}\text{ wins }\mathsf{Exp}^{\mathsf{NT}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)\right],$
is negligible. Here, $\mathsf{Exp}^{\mathsf{NT}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)$ denotes the $NT$ experiment between a challenger $\mathcal{C}$ and the adversary $\mathcal{A}$, as illustrated in Figure 2.
Setup: The challenger $\mathcal{C}$ begins by executing $\textsf{Setup}(1^{\lambda})$ to generate the public parameters $\mathsf{pp}$. It then generates key pairs for both the signer $\mathsf{S}$ and the designated verifier $\mathsf{V}$ by running $\textsf{SigKeyGen}(\mathsf{pp})$ and $\textsf{VerKeyGen}(\mathsf{pp})$, obtaining $(\mathsf{sk_{S}},\mathsf{pk_{S}})$ and $(\mathsf{sk_{V}},\mathsf{pk_{V}})$, respectively. The public keys $(\mathsf{pk_{S}},\mathsf{pk_{V}})$ are then provided to the adversary $\mathcal{A}$.
Challenge: The adversary $\mathcal{A}$ adaptively selects a target message $m^{\ast}$. In response, $\mathcal{C}$ generates a real signature $\sigma_{0}^{\ast}$ by running $\mathsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m^{\ast})$, and also computes a simulated signature $\sigma_{1}^{\ast}$ using $\mathsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m^{\ast})$. $\mathcal{C}$ chooses a random bit $b\in\{0,1\}$, and returns the signature $\sigma_{b}^{\ast}$ to $\mathcal{A}$.
Output: Eventually, $\mathcal{A}$ outputs a bit $b^{\ast}$. The challenge is considered successful if $b^{\ast}=b$.
Figure 2: $\mathsf{Exp}_{\mathsf{SDVS},\mathcal{A}}^{\mathsf{NT}}$ Experiment.
The advantage of $\mathcal{A}$ in the above game is defined by
$\mathsf{Adv}^{\mathsf{NT}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)=\left|\Pr[b^{\ast}=b]-\frac{1}{2}\right|,$
where the probability is taken over the randomness of $\mathcal{A}$, and the algorithms $\mathsf{Sign}$ and $\mathsf{Simul}$.
Privacy of Signer’s Identity (PSI) Security
The privacy of signer’s identity (PSI) security notion ensures that the identity of the actual signer remains hidden. Within an SDVS scheme, even when a signature and two valid signing key pairs are available, no adversary observing the communication between the signer and the designated verifier can determine—except with negligible probability—which of the two secret keys was used to generate the signature. This anonymity is achieved by incorporating the designated verifier’s secret key into the verification procedure, making it infeasible to link the signature to a specific signer.
Definition 2.4 (PSI Security).
A strong designated verifier signature (SDVS) scheme satisfies privacy of signer’s identity (PSI) security if the advantage of any PPT adversary $\mathcal{A}$, defined as
$\mathsf{Adv}^{\mathsf{PSI}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)=\Pr\left[\mathcal{A}\text{ wins }\mathsf{Exp}^{\mathsf{PSI}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)\right],$
is negligible. Here, $\mathsf{Exp}^{\mathsf{PSI}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)$ denotes the PSI experiment between a challenger $\mathcal{C}$ and the adversary $\mathcal{A}$, as illustrated in Figure 3.
Setup: The challenger $\mathcal{C}$ begins by executing $\textsf{Setup}(1^{\lambda})$ to generate the public parameters $\mathsf{pp}$. Next, it invokes $\textsf{SigKeyGen}(\mathsf{pp})$ twice to obtain two signing key pairs: $(\mathsf{sk_{S_{0}}},\mathsf{pk_{S_{0}}})$ and $(\mathsf{sk_{S_{1}}},\mathsf{pk_{S_{1}}})$, corresponding to signers $\mathsf{S_{0}}$ and $\mathsf{S_{1}}$, respectively. It also runs $\textsf{VerKeyGen}(\mathsf{pp})$ to generate the verifier’s key pair $(\mathsf{sk_{V}},\mathsf{pk_{V}})$. The challenger keeps $\mathsf{sk_{V}}$ private and sends the tuple $(\mathsf{pp},\mathsf{sk_{S_{0}}},\mathsf{pk_{S_{0}}},\mathsf{sk_{S_{1}}},\mathsf{pk_{S_{1}}},\mathsf{pk_{V}})$ to the adversary $\mathcal{A}$.
Query Phase: In this phase, $\mathcal{C}$ responds to polynomially many adaptive queries made by $\mathcal{A}$ by following the steps described below: Simulating Oracle $\mathcal{O}_{\mathsf{Simul}}$: Upon receiving queries on message $m$ and a bit $b$ from $\mathcal{A}$, $\mathcal{C}$ computes a simulated signature $\sigma_{b}$ by executing $\textsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S_{b}}},m)$ and returns it to $\mathcal{A}$. Verifying Oracle: $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving queries on message-signature pairs $(m,\sigma)$ along with a bit $b\in\{0,1\}$ from $\mathcal{A}$, $\mathcal{C}$ computes $\textsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S_{b}}},\sigma,m)$ and returns 1 if the signature is valid and 0 otherwise. Challenge: The adversary $\mathcal{A}$ selects a challenge message $m^{\ast}$. Upon this, $\mathcal{C}$ computes two genuine signatures: $\sigma_{0}^{\ast}$ by running $\textsf{Sign}(\mathsf{pp},\mathsf{sk_{S_{0}}},\mathsf{pk_{V}},m^{\ast})$, and $\sigma_{1}^{\ast}$ by running $\textsf{Sign}(\mathsf{pp},\mathsf{sk_{S_{1}}},\mathsf{pk_{V}},m^{\ast})$. Then, a bit $b\in\{0,1\}$ is chosen uniformly at random, and $\sigma^{\ast}_{b}$ is sent to $\mathcal{A}$.
Output: Eventually, $\mathcal{A}$ outputs a guess $b^{\ast}\in\{0,1\}$. It succeeds if both of the following conditions are met: $b^{\ast}=b$; The pair $(\sigma_{b^{\ast}}^{\ast},m^{\ast})$ was not previously submitted in a verification query for $b^{\ast}\in\{0,1\}$.
Figure 3: $\mathsf{Exp}_{\mathsf{SDVS},\mathcal{A}}^{\mathsf{PSI}}$ Experiment.
The advantage of $\mathcal{A}$ in the above game is defined by
$\mathsf{Adv}^{\mathsf{PSI}}_{\mathsf{SDVS},\mathcal{A}}(\lambda)=\left|\Pr[b^{\ast}=b]-\frac{1}{2}\right|,$
where the probability is taken over the randomness of $\mathcal{A}$, and the algorithms $\mathsf{Sign}$ and $\mathsf{Simul}$.
2.3 Elliptic Curves and Isogenies
This section recalls several fundamental properties of supersingular elliptic curves that are relevant to our construction. For a comprehensive treatment of the theory of elliptic curves, we refer the reader to [25, 26].
Let $k:=\mathbb{F}_{q}$ denote a finite field, where $q:=p^{n}$ for a prime $p>3$ and a positive integer $n$. An elliptic curve $E$ over $k$ is a smooth, projective, genus-1 curve defined over $k$ with a distinguished point $\mathcal{O}_{E}$ serving as the identity element for the group law on $E$. For a positive integer $\ell$, the $\ell$-torsion subgroup of $E$ is defined as $E[\ell]:=\{P\in E(\overline{k})\mid[\ell]P=\mathcal{O}_{E}\}$. An elliptic curve is called supersingular if its group of $p$-torsion points is trivial over $\overline{\mathbb{F}}_{p}$, i.e., $E[p]=\{\mathcal{O}_{E}\}$. In particular, if $E/\mathbb{F}_{p}$ is supersingular, then #$\#E(\mathbb{F}_{p})=p+1$.
An isogeny is a surjective morphism between elliptic curves that also preserves the group structure. Two elliptic curves $E_{1}$ and $E_{2}$ defined over $\mathbb{F}_{q}$ are said to be isogenous if there exists an isogeny between them. For any finite subgroup $G\subset E(\mathbb{F}_{q})$, there exists a unique (up to isomorphism) separable isogeny $\phi:E\to E^{\prime}:=E/G$ with kernel $\ker(\phi)=G$.
An endomorphism is an isogeny from an elliptic curve $E$ to itself. Examples include the multiplication-by-$m$ map $[m]:P\mapsto[m]P$ for $m\in\mathbb{Z}$, and the Frobenius endomorphism $\pi:(x,y)\mapsto(x^{q},y^{q})$ for an elliptic curve defined over $\mathbb{F}_{q}$. The set of all endomorphisms of $E$ forms a ring under addition and composition, known as the endomorphism ring of $E$ and denoted by $\mathrm{End}(E)$. For a supersingular elliptic curve $E$ defined over $\mathbb{F}_{p^{2}}$, the $\mathrm{End}(E)$ is isomorphic to a maximal order in a quaternion algebra that is ramified at $p$ and $\infty$. In contrast, for a supersingular elliptic curve defined over $\mathbb{F}_{p}$, the ring of $\mathbb{F}_{p}$-rational endomorphisms, denoted by $\mathrm{End}_{\mathbb{F}_{p}}(E)$, is isomorphic to an order in some imaginary quadratic field. Hence, we have a strict inclusion $\mathrm{End}_{\mathbb{F}_{p}}(E)\subset\mathrm{End}(E)$.
Let $\mathfrak{O}\subset\mathbb{Q}(\sqrt{-p})$ be an order. The ideal class group $\mathrm{Cl}(\mathfrak{O})$ is defined as the quotient $\mathcal{I}_{\mathfrak{O}}/\mathcal{P}_{\mathfrak{O}}$, where $\mathcal{I}_{\mathfrak{O}}$ denotes the group of invertible fractional ideals and $\mathcal{P}_{\mathfrak{O}}$ denotes the subgroup of principal fractional ideals. The class group $\mathrm{Cl}(\mathfrak{O})$ acts naturally on the set $\mathcal{E}ll_{p}(\mathfrak{O})$ of $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves defined over $\mathbb{F}_{p}$ with endomorphism ring isomorphic to a imaginary quadratic order $\mathfrak{O}$. For an $\mathfrak{O}$-ideal $\mathfrak{a}$, the associated isogeny is determined by the kernel subgroup
$S_{\mathfrak{a}}:=\bigcap_{\alpha\in\mathfrak{a}}\ker(\alpha),$
yielding the isogenous curve $\mathfrak{a}\ast E:=E/S_{\mathfrak{a}}$. The isogeny $\phi_{\mathfrak{a}}:E\rightarrow E/S_{\mathfrak{a}}$ is well-defined and unique up to $\mathbb{F}_{p}$-isomorphism. This gives rise to the group action
$\ast:\mathrm{Cl}(\mathfrak{O})\times\mathcal{E}ll_{p}(\mathfrak{O})\rightarrow\mathcal{E}ll_{p}(\mathfrak{O}),$
which is both free and transitive.
Notation. Following [24], if $\mathrm{Cl}(\mathfrak{O})$ is cyclic with generator $\mathfrak{g}$, then the map $\mathbb{Z}/N\mathbb{Z}\xhookrightarrow{}\mathrm{Cl}(\mathfrak{O})$ defined by $a\mapsto\mathfrak{g}^{a}$ gives a convenient representation of class group elements. Here, $N=\#\mathrm{Cl}(\mathfrak{O})$, and any element $\mathfrak{a}\in\mathrm{Cl}(\mathfrak{O})$ can be written as $\mathfrak{g}^{a}$ for some $a\in\mathbb{Z}/N\mathbb{Z}$. Using the shorthand $[a]$ for $\mathfrak{g}^{a}$ and $[a]E:=\mathfrak{g}^{a}*E$, it follows that $[a][b]E=[a+b]E$.
2.4 Hardness Assumptions
We begin by recalling the basic framework of one-way group actions and the notion of Hard Homogeneous Spaces (HHS), which form the foundation for isogeny-based hardness assumptions used in our construction.
Definition 2.5.
Given a group $G$ with identity element $e$, and a set $X$, then a (left) group action $\ast$ of $G$ on $X$ is a function
$\ast:G\times X\rightarrow X,$
satisfying the following axioms:
  • (i)
    Identity: $e\ast x=x$ for all $x\in X$.
  • (ii)
    Compatibility: $g\ast(h\ast x)=(gh)\ast x$ for all $g,h\in G$ and all $x\in X$.
To construct public-key cryptographic schemes based on group actions, Couveignes [27] introduced the concept of the HHS, defined as follows.
Hard Homogeneous Spaces (HHS).
A (principal) homogeneous space for a group $G$ is a non-empty finite set $X$ on which $G$ acts freely and transitively. That is, for every $x,y\in X$, there exists a unique $g\in G$ such that $y=g\ast x$. A hard homogeneous space (HHS) is a finite principal homogeneous space that satisfies the following two conditions:
i.
The following operations are efficiently computable (i.e., executable in polynomial time):
$\bullet$
(Group Operations) Given strings $g$ and $h$, decide if they represent elements in $G$ and whether they are equal. Given $g,h\in G$, compute $g^{-1}$, $gh$, and decide if $g=h$.
$\bullet$
(Random Element) Sample a random element in $G$ with uniform probability.
$\bullet$
(Membership) Given a string $x_{0}$, decide if $x_{0}$ represents an element in $X$.
$\bullet$
(Equality) Given $x_{1},x_{2}\in X$, decide if $x_{1}=x_{2}$.
$\bullet$
(Action) Given $g\in G$ and $x\in X$, then compute $g\ast x$.
ii.
The following computational problems are assumed to be hard (i.e., infeasible to solve in polynomial time):
$\bullet$
Vectorization: Given $x,y\in X$, find $g\in G$ such that $y=g\ast x$.
$\bullet$
Parallelization: Given $x,y,z\in X$ such that $y=g\ast x$, compute $z^{\prime}=g\ast z$.
Our isogeny-based construction, presented in Section 3, is built upon the hardness of the Group Action Inverse Problem (GAIP), which captures the difficulty of inverting group actions induced by ideal class groups on supersingular elliptic curves. This assumption underpins the security of the CSI-FiSh signature scheme [24].
Problem 2.6 (Group Action Inverse Problem (GAIP) [24]).
Given two supersingular elliptic curves $E$ and $E^{\prime}$ over the same finite field $\mathbb{F}_{p}$ with $\operatorname{End}_{\mathbb{F}_{p}}(E)\cong\operatorname{End}_{\mathbb{F}_{p}}(E^{\prime})\cong\mathfrak{O}$, the goal is to find an ideal $\mathfrak{a}\subset\mathfrak{O}$ such that $E^{\prime}=\mathfrak{a}*E$.
The security of CSI-FiSh—and by extension, the CSI-SDVS scheme introduced in Section 3—relies on a stronger assumption known as the Multi-Target Group Action Inverse Problem (MT-GAIP). As demonstrated in [28], MT-GAIP is tightly reducible to GAIP when the structure of the class group is known, which is the case in the CSIDH-based setting.
Problem 2.7 (Multi-Target Group Action Inverse Problem (MT-GAIP) [24]).
Let $E_{0},E_{1},\ldots,E_{k}$ be $k+1$ supersingular elliptic curves defined over $\mathbb{F}_{p}$, with $\operatorname{End}_{\mathbb{F}_{p}}(E_{i})\cong\mathfrak{O}$ for all $0\leq i\leq k$. Find an ideal $\mathfrak{a}\subset\mathfrak{O}$ such that $E_{i}=\mathfrak{a}*E_{j}$ for some distinct indices $i,j\in\{0,\ldots,k\}$.
The best known classical algorithm for solving GAIP (Problem 2.6) and its multi-target variant (Problem 2.7) runs in time $\mathcal{O}(\sqrt{N})$, where $N=\#\operatorname{Cl}(\mathfrak{O})$. On the quantum side, Kuperberg’s algorithm for the hidden shift problem [29, 30] achieves subexponential complexity. However, its concrete effectiveness in this setting remains an active topic of research [31, 32].
3 The Scheme: CSI-SDVS
In this section, we present CSI-SDVS, a new isogeny-based SDVS scheme. The construction integrates the CSIDH framework with the exact group action structure introduced in CSI-FiSh. This hybrid design preserves the post-quantum security guarantees of class group actions. By leveraging the explicitly known cyclic structure of the class group for CSIDH-friendly primes (as shown in [24]), deterministic computation and uniform randomness are naturally enabled. As a result, we achieve efficient and verifiable designated signatures without relying on heuristic assumptions.
In our scheme, CSI-SDVS, the key generation procedures for both the signer and the designated verifier closely follow the key generation phase in CSI-FiSh. At the heart of the $\mathsf{Sign}$ procedure—also mirrored in $\mathsf{Simul}$—lies a structured interaction between class group elements and $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves. Starting from a base curve $E_{0}$, the signer selects a secret randomness $b\in\mathbb{Z}_{N}$ and computes the intermediate curve $\bar{E}=[b]E_{0}$ using the known group action. The designated verifier, holding a secret key $v\in\mathbb{Z}_{N}$, then computes the target curve $Y=[v]\bar{E}=[v+b]E_{0}$, which serves as the core component for verifying the designated verifier signature.
The complete scheme consists of six polynomial-time algorithms, described in detail in the following procedures.
$\mathsf{CSI\text{-}SDVS.Setup}(1^{\lambda})\rightarrow\mathsf{pp}$:
A trusted party executes the setup procedure to generate the public parameters $\mathsf{pp}$ as follows. First, it selects a large prime $p=4\ell_{1}\ell_{2}\dots\ell_{n}-1$, where the $\ell_{i}$ are small, distinct, odd primes. Specifically, it sets $n=74$, $\ell_{1}=3$, $\ell_{73}=373$, and $\ell_{74}=587$. Next, it fixes a base curve $E_{0}:y^{2}=x^{3}+x$ defined over $\mathbb{F}_{p}$, such that $E_{0}\in\mathcal{E}ll_{p}(\mathfrak{O})$, and chooses a generator $\mathfrak{g}$ of the class group $\mathcal{G}=\text{Cl}(\mathfrak{O})$ with class number $N\approx 2^{\lambda}\approx p^{1/2}$, where $\mathfrak{O}=\mathbb{Z}[\sqrt{-p}]$. It then samples a cryptographically secure hash function :$\mathcal{H}:\{0,1\}^{\ast}\rightarrow\{0,1\}^{\lambda}$, where $\lambda$ is a security parameter. In addition, a positive integer $\eta$—which depends on the security parameter—is specified. Finally, it outputs the public parameters $\mathsf{pp}:=(p,\mathfrak{g},N,E_{0},\mathcal{H},\eta)$.
$\mathsf{CSI\text{-}SDVS.SigKeyGen}(\mathsf{pp})\rightarrow(\mathsf{sk_{S}},\mathsf{pk_{S}})$:
Given $\mathsf{pp}$, the signer generates a key pair as follows:
  1. 1.
    Sample $s_{i}\in\mathbb{Z}_{N}$ uniformly at random and compute $E_{i}:=[s_{i}]E_{0}$ for each $i\in[\eta]$.
  2. 2.
    Set the signer’s secret/public key pair as $(\mathsf{sk_{S}},\mathsf{pk_{S}}):=(\{s_{i}\}_{i=1}^{\eta},\{E_{i}\}_{i=1}^{\eta})$.
$\mathsf{CSI\text{-}SDVS.VerKeyGen}(\mathsf{pp})\rightarrow(\mathsf{sk_{V}},\mathsf{pk_{V}})$:
Given $\mathsf{pp}$, the designated verifier generates a key pair as follows:
  1. 1.
    Sample $v_{i}\in\mathbb{Z}_{N}$ uniformly at random and compute $\hat{E}_{i}:=[v_{i}]E_{0}$ for each $i\in[\eta]$.
  2. 2.
    Set the verifier’s secret/public key pair as $(\mathsf{sk_{V}},\mathsf{pk_{V}}):=(\{v_{i}\}_{i=1}^{\eta},\{\hat{E}_{i}\}_{i=1}^{\eta})$.
$\mathsf{CSI\text{-}SDVS.Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m)\rightarrow\sigma$:
Employing $\mathsf{pp}$, the secret key $\mathsf{sk_{S}}=\{s_{i}\}_{i=1}^{\eta}$, and the designated verifier’s public key $\mathsf{pk_{V}}=\{\hat{E}_{i}\}_{i=1}^{\eta}$, a signer computes the designated verifier signature $\sigma$ on $m\in\{0,1\}^{\ast}$ as follows:
  1. 1.
    Sample $b_{i}\in\mathbb{Z}_{N}$ uniformly at random for each $i\in[\eta]$.
  2. 2.
    Compute the curves $Y_{i}:=[b_{i}]\hat{E}_{i}$ for each $i\in[\eta]$.
  3. 3.
    Compute the hash value $h:=\mathcal{H}(Y_{1}\|\ldots\|Y_{\eta}\|m)$.
  4. 4.
    Compute the vector $z_{i}:=b_{i}-s_{i}\mod N$ for each $i\in[\eta]$.
  5. 5.
    Output the designated verifier signature $\sigma:=(h,\mathbf{z})$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$.
$\mathsf{CSI\text{-}SDVS.Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m,\sigma)\rightarrow\{0,1\}$:
Incorporating $\mathsf{pp}$, the secret key $\mathsf{sk_{V}}=\{v_{i}\}_{i=1}^{\eta}$, and the signer’s public key $\mathsf{pk_{S}}=\{E_{i}\}_{i=1}^{\eta}$, a designated verifier checks the validity of the designated verifier signature $\sigma$ on $m\in\{0,1\}^{\ast}$ as follows:
  1. 1.
    Parse $\sigma:=(h,\mathbf{z})$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$.
  2. 2.
    Compute $y_{i}:=v_{i}+z_{i}\mod N$ for each $i\in[\eta]$.
  3. 3.
    Compute the curves $Y^{\prime}_{i}=[y_{i}]E_{i}$ for each $i\in[\eta]$.
  4. 4.
    Recover the hash value $h^{\prime}:=\mathcal{H}(Y^{\prime}_{1}\|\ldots\|Y^{\prime}_{\eta}\|m)$.
  5. 5.
    Output 1 if $h^{\prime}=h$, otherwise 0.
$\mathsf{CSI\text{-}SDVS.Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m)\rightarrow\sigma$:
Using $\mathsf{pp}$, the secret key $\mathsf{sk_{V}}=\{v_{i}\}_{i=1}^{\eta}$, and the signer’s public key $\mathsf{pk_{S}}=\{E_{i}\}_{i=1}^{\eta}$, a designated verifier computes the simulated designated verifier signature $\sigma$ on $m\in\{0,1\}^{\ast}$ as follows:
  1. 1.
    Sample $r_{i}\in\mathbb{Z}_{N}$ uniformly at random for each $i\in[\eta]$.
  2. 2.
    Compute the curves $Y_{i}:=[r_{i}]E_{i}$ for each $i\in[\eta]$.
  3. 3.
    Compute the hash value $h:=\mathcal{H}(Y_{1}\|\ldots\|Y_{\eta}\|m)$.
  4. 4.
    Compute the vector $z_{i}:=r_{i}-v_{i}\mod N$ for each $i\in[\eta]$.
  5. 5.
    Output the simulated designated verifier signature $\sigma:=(h,\mathbf{z})$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$.
Correctness.
The proposed scheme satisfies the correctness requirement stated in Definition 2.1. In particular, during the verification phase, the designated verifier validates a message-signature pair $(m,\sigma)$ using the inputs $(\mathsf{pp},\mathsf{sk}_{\mathsf{V}},\mathsf{pk}_{\mathsf{S}})$, where $\mathsf{sk_{V}}=\{v_{i}\}_{i=1}^{\eta}$ and $\mathsf{pk_{S}}=\{E_{i}\}_{i=1}^{\eta}$. We distinguish two correctness cases corresponding to the two ways a designated verifier signature may be generated, and verify that in both cases the verification algorithm accepts.
i.
When the designated verifier signature is generated by the actual signer: Let $\sigma:=(h,\mathbf{z})$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$. The verifier computes each $Y_{i}=[b_{i}]\hat{E}_{i}$ as follows:
$Y_{i}$ $=[v_{i}][z_{i}]E_{i}$
$=[v_{i}+b_{i}-s_{i}][s_{i}]E_{0}$
$=[v_{i}+b_{i}]E_{0}$
$=[b_{i}]\hat{E}_{i},$
which implies that $h=\mathcal{H}(Y_{1}\|\ldots\|Y_{\eta}\|m)$, as expected.
ii.
When the signature is simulated by the designated verifier: Let $\sigma^{\prime}:=(h,\mathbf{z})$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$ is computed using the Simul algorithm. Then each value $Y_{i}=[r_{i}]E_{i}$ is recovered by computing:
$Y_{i}$ $=[v_{i}][z_{i}]E_{i}$
$=[v_{i}+r_{i}-v_{i}]E_{i}$
$=[r_{i}]E_{i},$
which again yields $h=\mathcal{H}(Y_{1}\|\ldots\|Y_{\eta}\|m)$.
4 Efficiency Analysis
In this section, we analyze the communication and storage overhead of the proposed CSI-SDVS scheme.
The secret keys $\mathsf{sk}_{\mathsf{S}}=\{s_{i}\bmod N\}_{i=1}^{\eta}$ and $\mathsf{sk}_{\mathsf{V}}=\{v_{i}\bmod N\}_{i=1}^{\eta}$, held by the signer and the designated verifier respectively, consist of $\eta$ elements in $\mathbb{Z}_{N}$. Since $\log(N)=\lambda$, the total size of each secret key is $\eta\cdot\lambda=\mathcal{O}(\lambda)$ bits.
Each curve in $\mathcal{E}ll_{p}(\mathfrak{O})$ is uniquely represented by a Montgomery coefficient $A\in\mathbb{F}_{p}$, defining a curve $E_{A}:y^{2}=x^{3}+Ax^{2}+x$. Hence, the public keys $\mathsf{pk}_{\mathsf{S}}$ and $\mathsf{pk}_{\mathsf{V}}$ each contain $\eta$ such elements over $\mathbb{F}_{p}$, yielding a size of $\eta\log(p)=\mathcal{O}(\lambda)$ bits.
A signature $\sigma=(h,\mathbf{z})$, where $\mathbf{z}=\{z_{i}\bmod N\}_{i=1}^{\eta}$, consists of a $\lambda$-bit hash and $\eta$ elements in $\mathbb{Z}_{N}$, each of $\lambda$ bits. This results in a total signature size of $\lambda(1+\eta)=\mathcal{O}(\lambda)$ bits.
Following CSI‑FiSh and CSI‑SharK [33], we adopt $\eta=16$ for 128-bit security, striking a balance between soundness error, key size, and efficiency. Higher values such as $\eta=32$ may be chosen to further reduce the soundness error at the expense of increased overhead.
In comparison, lattice-based SDVS schemes [14, 15, 16, 17] typically require key and signature sizes of $\mathcal{O}(\lambda^{2})$ due to high-dimensional trapdoor structures. Our isogeny-based construction avoids such overhead by relying on group actions in fixed-size cyclic groups, achieving compact $\mathcal{O}(\lambda)$ complexity across all components.
5 Security Analysis of CSI-SDVS
Theorem 5.1.
The CSI-SDVS scheme introduced in Section 3 is strongly unforgeable under chosen message attacks (SUF-CMA), as defined in Definition 2.2.
Proof.
We prove that the CSI-SDVS scheme satisfies SUF-CMA security under Definition 2.2. Suppose, for the sake of contradiction, that there exists a PPT adversary $\mathcal{A}$ that succeeds in the experiment $\mathsf{Exp}^{\mathsf{SUF\text{-}CMA}}_{\mathsf{CSI\text{-}SDVS},\mathcal{A}}(\lambda)$ (illustrated in Figure 1) with non-negligible probability. Acting as the challenger, a simulator $\mathcal{S}$ interacts with $\mathcal{A}$ by responding to its queries through the signing oracle $\mathcal{O}_{\mathsf{Sign}}$, the simulating oracle $\mathcal{O}_{\mathsf{Simul}}$, the verification oracle $\mathcal{O}_{\mathsf{Verify}}$, and the random oracle $\mathcal{O}_{\mathcal{H}}$. At the end of the interaction, $\mathcal{A}$ outputs a forged signature $\sigma^{\ast}:=(h^{\ast},\mathbf{z}^{\ast})$ on a message $m^{\ast}$ with non-negligible probability, such that
$Y^{\ast}_{i}=[v_{i}+z^{\ast}_{i}]E_{i},\quad\text{and}\quad h^{\ast}=\mathcal{H}(Y^{\ast}_{1}\|\dots\|Y^{\ast}_{\eta}\|m^{\ast})\quad\text{for all }i\in[\eta].$
Leveraging the forgery $\sigma^{\ast}$, the simulator $\mathcal{S}$ computes the secret curves $[s_{i}+v_{i}]E_{0}$ from the public values $[s_{i}]E_{0}$ and $[v_{i}]E_{0}$. This contradicts the assumed hardness of the parallelization problem in the HHS model, as defined in Section 2.4.
Setup. Given public parameters $\mathsf{pp}:=(p,\mathfrak{g},N,E_{0},\mathcal{H},\eta)$, as defined in Section 3, the simulator $\mathcal{S}$ proceeds as follows.
  1. 1.
    For each $i\in[\eta]$, sample $s_{i}\in\mathbb{Z}_{N}$ uniformly at random and compute $E_{i}:=[s_{i}]E_{0}$. Set the signer’s key pair as $(\mathsf{sk_{S}},\mathsf{pk_{S}}):=(\{s_{i}\}_{i=1}^{\eta},\{E_{i}\}_{i=1}^{\eta})$.
  2. 2.
    For each $i\in[\eta]$, sample $v_{i}\in\mathbb{Z}_{N}$ uniformly at random and compute $\hat{E}_{i}:=[v_{i}]E_{0}$. Set the verifier’s key pair as $(\mathsf{sk_{V}},\mathsf{pk_{V}}):=(\{v_{i}\}_{i=1}^{\eta},\{\hat{E}_{i}\}_{i=1}^{\eta})$.
  3. 3.
    Send $(\mathsf{pp},\mathsf{pk_{S}},\mathsf{pk_{V}})$ to the adversary $\mathcal{A}$, while keeping $(\mathsf{sk_{S}},\mathsf{sk_{V}})=(\{s_{i}\}_{i=1}^{\eta},\{v_{i}\}_{i=1}^{\eta})$ secret, and define sets $\mathcal{Q}:=\emptyset$, $\mathsf{HList}:=\emptyset$.
Simulation of Queries. $\mathcal{A}$ issues polynomially many adaptive queries to the oracles $\mathcal{O}_{\mathcal{H}}$, $\mathcal{O}_{\mathsf{Sign}}$, $\mathcal{O}_{\mathsf{Simul}}$, and $\mathcal{O}_{\mathsf{Verify}}$.
  • Hashing Oracle $\mathcal{O}_{\mathcal{H}}$: Upon receiving a query on a tuple $(m,\{Y_{i}\}_{i=1}^{\eta})$, the simulator $\mathcal{S}$ checks whether a corresponding hash output $h$ is already stored in the hash list $\mathsf{HList}$. If such an entry exists, $\mathcal{S}$ returns $h$ directly. Otherwise, it samples $h\xleftarrow{\$}\{0,1\}^{\lambda}$ at random, programs the oracle as $h=:\mathcal{H}(Y_{1}\|\ldots\|Y_{\eta}\|m)$ and returns $h$ to $\mathcal{A}$. Finally, $\mathcal{S}$ updates the hash list as $\mathsf{HList}\leftarrow\mathsf{HList}\cup\{(m,\{Y_{i}\}_{i=1}^{\eta},h,\perp)\}$.
  • Signing Oracle $\mathcal{O}_{\mathsf{Sign}}$: In response to a query on a message $m$, $\mathcal{S}$ produces a signature as follows.
    • Sample $z_{i}\in\mathbb{Z}_{N}$, and compute the curves $\bar{E}_{i}:=[z_{i}]E_{i}$ for each $i\in[\eta]$.
    • Sample $v_{i}\in\mathbb{Z}_{N}$, and compute the curves $Y_{i}:=[v_{i}]\bar{E}_{i}$ for each $i\in[\eta]$.
    • Sample $h\xleftarrow{\$}\{0,1\}^{\lambda}$ at random, program the oracle as $h=:\mathcal{H}(Y_{1}\|\ldots\|Y_{\eta}\|m)$, and update the hash list as $\mathsf{HList}\leftarrow\mathsf{HList}\cup\{(m,\{Y_{i}\}_{i=1}^{\eta},h,\{\bar{E}_{i}\}_{i=1}^{\eta})\}$.
    • Return $\sigma:=(h,\mathbf{z})$ to $\mathcal{A}$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$. It updates the set $\mathcal{Q}$ as $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$.
  • Simulating Oracle $\mathcal{O}_{\mathsf{Simul}}$: In response to a query on a message $m$, the simulator $\mathcal{S}$ produces a simulated designated verifier signature with a similar arguments as in the Signing Oracle $\mathcal{O}_{\mathsf{Sign}}$, updates the set $\mathcal{Q}$ as $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(m,\sigma)\}$, and returns a signature $\sigma:=(h,\mathbf{z})$ to $\mathcal{A}$.
  • Verifying oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving a query on message-signature pair $(m,\sigma)$ to this oracle, $\mathcal{S}$ proceeds as follows.
    • Parse $\sigma:=(h,\mathbf{z})$, where $\mathbf{z}:=\{z_{i}\}_{i=1}^{\eta}$.
    • Check whether there exist an entry $(m,\{Y_{i}\}_{i=1}^{\eta},h,\{\bar{E}_{i}\}_{i=1}^{\eta})\in\mathsf{HList}$ such that $m=m$ and $h=h$
    • Extract $\{\bar{E}_{i}\}_{i=1}^{\eta}$ from $(m,\{Y_{i}\}_{i=1}^{\eta},h,\{\bar{E}_{i}\}_{i=1}^{\eta})$, and check if $[z_{i}]E_{i}=\bar{E}_{i}$ or for all $i\in[\eta]$.
    • Output 1 if all the above are satisfied, otherwise 0.
Extracting the Forgery. Eventually, the adversary $\mathcal{A}$ outputs a forgery $(m^{\ast},\sigma^{\ast})$ with non-negligible probability, where $\sigma^{\ast}:=(h^{\ast},\mathbf{z}^{\ast})$. The simulator $\mathcal{S}$ then searches for a matching entry $(m,h,\{Y_{i}\}_{i=1}^{\eta},\perp)$ in the hash list $\mathsf{HList}$ such that $m^{\ast}=m$, $h^{\ast}=h$. If such an entry is found, $\mathcal{S}$ can solve an instance of the parallelization problem in the HHS model, as given in Section 2.4, with a non-negligible probability by recovering the secret value $[s_{i}+v_{i}]E_{0}$ as follows:
$[-z^{\ast}_{i}]Y_{i}=[s_{i}-b^{\ast}_{i}][b^{\ast}_{i}]\hat{E}_{i}=[s_{i}]\hat{E}_{i}=[s_{i}+v_{i}]E_{0},\quad\text{for each }i\in[\eta],$
where $\{s_{i}\}_{i=1}^{\eta}$ and $\{v_{i}\}_{i=1}^{\eta}$ are the secret keys of the signer and verifier, respectively. ∎
Theorem 5.2.
The CSI-SDVS scheme presented in Section 3 satisfies non-transferability (NT) security, as per Definition 2.3.
Proof.
According to the NT security model of the SDVS scheme defined in Definition 2.3, a game is established between the PPT distinguisher $\mathcal{A}$ and the simulator $\mathcal{B}$ as follows.
Setup: $\mathcal{B}$ performs the following operations:
  • 1.
    Execute $\mathsf{Setup}(1^{\lambda})$ to obtain the public parameters $\mathsf{pp}$.
  • 2.
    Run $\mathsf{SigKeyGen}(\mathsf{pp})$ to generate the signer’s key pair $(\mathsf{sk_{S}},\mathsf{pk_{S}})=(\{s_{i}\}_{i=1}^{\eta},\{E_{i}\}_{i=1}^{\eta})$.
  • 3.
    Run $\mathsf{VerKeyGen}(\mathsf{pp})$ to generate the designated verifier’s key pair $(\mathsf{sk_{V}},\mathsf{pk_{V}})=(\{v_{i}\}_{i=1}^{\eta},\{\hat{E}_{i}\}_{i=1}^{\eta})$.
  • 4.
    Provide $\mathcal{A}$ with $(\mathsf{pp},\mathsf{pk_{S}},\mathsf{pk_{V}})$, while retaining $\mathsf{sk_{S}}$ and $\mathsf{sk_{V}}$ as private.
Challenge: Upon receiving a message $m$ from $\mathcal{A}$, $\mathcal{B}$ proceeds as follows:
  • 1.
    Invoke $\mathsf{Sign}(\mathsf{pp},\mathsf{sk_{S}},\mathsf{pk_{V}},m)$ to obtain a real designated verifier signature $\sigma_{0}=(h^{(0)},\mathbf{z}^{(0)})$.
  • 2.
    Invoke $\mathsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S}},m)$ to obtain a simulated designated verifier signature $\sigma_{1}=(h^{(1)},\mathbf{z}^{(1)})$.
  • 3.
    Select a random bit $b\xleftarrow{\$}\{0,1\}$ and send the pair $(m,\sigma_{b})$ to $\mathcal{A}$.
Output: $\mathcal{A}$ responds with a guess $b^{\ast}\in\{0,1\}$.
In the above process, the randomness of the vectors $h^{(0)}$ and $h^{(1)}$ within the range $\{0,1\}^{\lambda}$ is ensured by the property of the hash function $\mathcal{H}$. For each $i=1,\ldots,\eta$, since the values $b^{(0)}_{i}$ and $b^{(1)}_{i}$ are sampled uniformly at random from $\mathbb{Z}_{N}$, then:
  • The components
    $z^{(0)}_{i}=b^{(0)}_{i}-s_{i}\mod N\quad\text{and}\quad z^{(1)}_{i}=b^{(1)}_{i}-v_{i}\mod N$
    retain their uniform randomness. Consequently, the vectors $\mathbf{z}^{(0)}=(z^{(0)}_{1},\ldots,z^{(0)}_{\eta})$ and $\mathbf{z}^{(1)}=(z^{(1)}_{1},\ldots,z^{(1)}_{\eta})$ are both uniformly distributed over $\mathbb{Z}_{N}^{\eta}$ and are thus identically distributed.
  • For the curves
    $[z^{(0)}_{i}]E_{i}=[b^{(0)}_{i}-s_{i}]E_{i}=[b^{(0)}_{i}]E_{0}\quad\text{and}\quad[z^{(1)}_{i}]E_{i}=[b^{(1)}_{i}-v_{i}]E_{i}=[b^{(1)}_{i}-v_{i}+s_{i}]E_{0},$
    the components
    $b^{(0)}_{i}\mod N\quad\text{and}\quad b^{(1)}_{i}-v_{i}+s_{i}\mod N$
    are uniformly distributed over $\mathbb{Z}_{N}$. Therefore, $[b^{(0)}_{i}]E_{0}$ and $[b^{(1)}_{i}-v_{i}+s_{i}]E_{0}$ are both uniformly distributed over $\mathcal{E}ll_{p}(\mathfrak{O})$.
As a result, the adversary $\mathcal{A}$ cannot distinguish whether the provided signature $\sigma_{b}$ was generated by the signer or simulated. Hence, its advantage in guessing $b$ correctly is negligible, thereby satisfying non-transferability as defined in Definition 2.3. ∎
Theorem 5.3.
If the MT-GAIP defined in Section 2.4 is hard, then the CSI-SDVS scheme introduced in Section 3 satisfies privacy of signer’s identity (PSI) security as defined in Definition 2.4.
Proof.
As specified in the PSI security model of the SDVS scheme, as per Definition 2.4, the interaction between the PPT adversary $\mathcal{A}$ and the challenger $\mathcal{C}$ is modeled as the following game.
Setup: The challenger $\mathcal{C}$ proceeds with the following steps:
  • 1.
    Execute $\mathsf{Setup}(1^{\lambda})$ to generate the public parameters $\mathsf{pp}$.
  • 2.
    Run $\mathsf{SigKeyGen}(\mathsf{pp})$ twice to obtain key pairs for signers $\mathsf{S_{0}}$ and $\mathsf{S_{1}}$ as follows:
    $(\mathsf{sk_{S_{0}}},\mathsf{pk_{S_{0}}})=(\{s^{(0)}_{i}\}_{i=1}^{\eta},\{E^{(0)}_{i}\}_{i=1}^{\eta}),\quad(\mathsf{sk_{S_{1}}},\mathsf{pk_{S_{1}}})=(\{s^{(1)}_{i}\}_{i=1}^{\eta},\{E^{(1)}_{i}\}_{i=1}^{\eta}).$
  • 3.
    Run $\mathsf{VerKeyGen}(\mathsf{pp})$ to generate the designated verifier’s key pair $(\mathsf{sk_{V}},\mathsf{pk_{V}})=(\{v_{i}\}_{i=1}^{\eta},\{\hat{E}_{i}\}_{i=1}^{\eta})$.
  • 4.
    Send $(\mathsf{pp},\mathsf{sk_{S_{0}}},\mathsf{pk_{S_{0}}},\mathsf{sk_{S_{1}}},\mathsf{pk_{S_{1}}},\mathsf{pk_{V}})$ to $\mathcal{A}$, while keeping $\mathsf{sk_{V}}$ secret.
Query Phase: During this phase, $\mathcal{C}$ responds to polynomially many adaptive queries from $\mathcal{A}$ as follows:
  • Simulation Oracle $\mathcal{O}_{\mathsf{Simul}}$: Upon receiving queries on message $m$ and a bit $b\in\{0,1\}$ from $\mathcal{A}$, $\mathcal{C}$ computes a simulated designated verifier signature by running: $\mathsf{Simul}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S_{b}}},m)$ and returns $\sigma_{b}=(h^{(b)},\mathbf{z}^{(b)})$ to $\mathcal{A}$.
  • Verification Oracle $\mathcal{O}_{\mathsf{Verify}}$: Upon receiving queries on message-signature pair $(m,\sigma)$ and bit $b\in\{0,1\}$ from $\mathcal{A}$, $\mathcal{C}$ runs: $\mathsf{Verify}(\mathsf{pp},\mathsf{sk_{V}},\mathsf{pk_{S_{b}}},\sigma,m)$ and returns $1$ if the signature is valid and 00 otherwise.
Challenge: Upon receiving a challenge message $m$ from $\mathcal{A}$, $\mathcal{C}$ executes the following:
  • 1.
    Compute a real designated verifier signature $\sigma_{0}$ from $\mathsf{S_{0}}$ by running:
    $\sigma_{0}=(h^{(0)},\mathbf{z}^{(0)})\xleftarrow{\$}\mathsf{Sign}(\mathsf{pp},\mathsf{sk_{S_{0}}},\mathsf{pk_{V}},m).$
  • 2.
    Similarly, compute a real designated verifier signature $\sigma_{1}$ from $\mathsf{S_{1}}$ by running:
    $\sigma_{1}=(h^{(1)},\mathbf{z}^{(1)})\xleftarrow{\$}\mathsf{Sign}(\mathsf{pp},\mathsf{sk_{S_{1}}},\mathsf{pk_{V}},m).$
  • 3.
    Choose a random bit $b\xleftarrow{\$}\{0,1\}$ and send $(m,\sigma_{b})$ to $\mathcal{A}$.
Output: Finally, $\mathcal{A}$ returns a guess bit $b^{\ast}\in\{0,1\}$.
Similar to the reasoning used in the proof of NT security in Theorem 5.2, the randomness of $h^{(0)}$ and $h^{(1)}$ is ensured by the properties of the hash function $\mathcal{H}$, with each vector uniformly distributed over $\{0,1\}^{\lambda}$. Likewise, the vectors $\mathbf{z}^{(0)}=\{z^{(0)}_{i}\}_{i=1}^{\eta}$ and $\mathbf{z}^{(1)}=\{z^{(1)}_{i}\}_{i=1}^{\eta}$ follow the same distribution as elements from $\mathbb{Z}_{N}^{\eta}$. Given a real message-signature pair $(m,\sigma_{b})$, the adversary $\mathcal{A}$ is expected to proceed as follows:
  1. 1.
    Parse $\sigma_{b}=(h^{(b)},\mathbf{z}^{(b)})=(h^{(b)}_{i},\{z^{(b)}_{i}\}_{i=1}^{\eta})$.
  2. 2.
    Compute $X^{(j)}_{i}:=[z^{(b)}_{i}]E^{(j)}_{i}$ for each $i\in[\eta]$ and $j=0,1$.
  3. 3.
    Compute $Y^{(j)}_{i}:=[v_{i}]X^{(j)}_{i}$ for each $i\in[\eta]$ and $j=0,1$.
  4. 4.
    Checck if $h^{(b)}=\mathcal{H}(Y^{(j)}_{1}\|\ldots\|Y^{(j)}_{\eta}\|m)$ for $j=0,1$.
According to the definition of MT-GAIP in Problem 2.7, it is evident that, without access to the designated verifier’s secret key $\mathsf{sk_{V}}=\{v_{i}\}_{i=1}^{\eta}$, even when $j=b$, the adversary $\mathcal{A}$ cannot compute the following set of curves:
$Y^{(j)}_{i}:=[v_{i}]X^{(j)}_{i},\quad\text{for all }i\in[\eta].$
in Step 3. As a result, the adversary cannot determine which signer produced the signature, as doing so would require recovering the curves $Y^{(b)}_{i}$ in order to compute $[s^{(b)}_{i}+v_{i}]E_{0}$:
$[-z^{(b)}_{i}][Y^{(b)}_{i}]$ $=[s^{(b)}_{i}-b^{(b)}_{i}][b^{(b)}_{i}]\hat{E}_{i}$
$=[s^{(b)}_{i}]\hat{E}_{i}$
$=[s^{(b)}_{i}+v_{i}]E_{0},$
which would enable distinguishing which of the two potential signers generated the signature $\sigma_{b}$. This guarantees that the signer’s anonymity is preserved, and that adversary’s distinguishing advantage remains negligible in the absence of $\mathsf{sk_{V}}=\{v_{i}\}_{i=1}^{\eta}$. ∎
Conclusion
In this work, we have introduced CSI-SDVS, a novel post-quantum Strong Designated Verifier Signature (SDVS) scheme founded on the class group action over the $\mathbb{F}_{p}$-isomorphism classes of supersingular elliptic curves, inspired by CSIDH and CSI-FiSh. Our construction achieves a compelling balance between security and efficiency, offering strong cryptographic guarantees such as Strong Unforgeability under Chosen-Message Attacks (SUF-CMA), Non-Transferability (NT), and Privacy of Signer’s Identity (PSI). One of the notable advantages of CSI-SDVS lies in its asymptotic efficiency: it achieves key and signature sizes of complexity $\mathcal{O}(\lambda)$, marking a significant improvement over the conventional $\mathcal{O}(\lambda^{2})$ bounds typical of lattice-based SDVS constructions. Although an earlier isogeny-based SDVS scheme [18] was proposed, it has been rendered insecure due to devastating cryptanalytic attacks [20, 21, 22] on SIDH. To the best of our knowledge, CSI-SDVS is currently the only isogeny-based SDVS construction that remains secure against all known quantum and classical attacks while offering both compactness and efficiency.